OpenStudy (anonymous):

Find the sum of all the numbers that can be formed with the digits 2,3,4,5 taken all at a time!!

OpenStudy (anonymous):

@Everyone

OpenStudy (anonymous):

Count repeats?

OpenStudy (anonymous):

@Ashleyisakitty @april115 @e.cociuba @thomaster @thomaster @Loujoelou @countonme123 @Muddingirl @MathFan12

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

So, 5555 would be a number

OpenStudy (anonymous):

@dimensionx @skullpatrol @satellite73

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

@KingGeorge @karatechopper

OpenStudy (kinggeorge):

First though is to write it as \[\sum_{i=2}^5\sum_{j=2}^5\sum_{k=2}^5\sum_{l=2}^5 (i\cdot 10^3+j\cdot10^2+k\cdot10+l)\]

OpenStudy (anonymous):

kay

OpenStudy (anonymous):

@KingGeorge Can u refer any good websites for permutations and combinations...........M kinda new to the topic!!

OpenStudy (kinggeorge):

I don't know of any good websites off the top of my head. But try searching for stuff about "combinatorics." This might get you some stuff you're interested in as long as you filter out some of the more advanced material.

OpenStudy (kinggeorge):

I don't know of any good ways to simplify the sum I wrote, but any decent graphing calculator or wolfram alpha should be able to sum that. I'll keep thinking to see if it can be simplified at all though.

OpenStudy (anonymous):

ta

OpenStudy (kinggeorge):

Are you sure that you can repeat digits?

OpenStudy (kinggeorge):

In any case, I did find a better solution. One minute while I type up a clear explanation.

OpenStudy (anonymous):

yes.sure........think I got it............The ans is 93324??

OpenStudy (kinggeorge):

Out of all of the 4-digit numbers you make, each one will have an equal probability of have a 2 in the first place, as it will a 5 or a 3 or a 4. And this is the same for the second/third/fourth places as well. So we can look at the average number that will be in these places. As it turns out, this will be \((2+3+4+5)/4=14/4=7/2\). Now, exactly how many numbers will there be total? We have 4 choices for each place, and 4 places. Thus, there are \(4^4=256\) different possibilities. So if we look at only a specific digit, the sum will be \(256*(7/2)=896\). Finally, we look at the different places. This results in the formula\[896*1000+896*100+896*10+896\]Run that through a calculator, and you should have an answer. And no, it is not 93324.

OpenStudy (anonymous):

I think Its 4! ways * 7/2 ..not 4^4

OpenStudy (kinggeorge):

We're allowing repetition, so it's 4^4. If repetition is not allowed, then yes, it would be 4!

OpenStudy (anonymous):

permutation so \[4P _{4}=4!\]

OpenStudy (kinggeorge):

Check to make absolute sure that you're including repetitions. Can you type out the exact wording here? Or did you already do that at the top?

OpenStudy (anonymous):

http://answers.yahoo.com/question/index?qid=20091115020541AA4ww3V I alredy did dat.......I found this and this matches with 93324

OpenStudy (kinggeorge):

That problem specifies that there are no repeats (that's what the "by permutation" means).

OpenStudy (anonymous):

I mentioned 4 comments ago,its permutation so 4!

OpenStudy (kinggeorge):

If you are indeed instructed to use permutations, then yes, the solution in that link is correct.

OpenStudy (anonymous):

so help me with this if it was not mentioned as permutations!!,we always shud take repetition into account??

OpenStudy (kinggeorge):

Correct. If the problem is as it was stated at the very top, then you should consider repetitions.

OpenStudy (anonymous):

But it is the same I typed!!!,given under permutation section!!

OpenStudy (kinggeorge):

Can you do me a favor and please type the problem EXACTLY as it was written?

OpenStudy (anonymous):

That is the xact question!!!

OpenStudy (kinggeorge):

"Find the sum of all the numbers that can be formed with the digits 2,3,4,5 taken all at a time?"

OpenStudy (anonymous):

yes..........King!!

OpenStudy (kinggeorge):

If that's the exact phrasing, then I think you need to consider repetitions.

OpenStudy (anonymous):

O'cmon..........Even if its given under permutation section,need I consider repetitions??

OpenStudy (anonymous):

dats a gen question,btw!!

OpenStudy (kinggeorge):

If it's in the permutation section, then maybe you should use permutations, in which case repetitions wouldn't be considered. Go ahead and use the permutations then. At worst, you should still get most of the credit for having the right method.

OpenStudy (anonymous):

guess wat.......Its 1 point question!!

OpenStudy (kinggeorge):

Go with the permutations then.

OpenStudy (anonymous):

And to think I just hav to waste dat long time for a 1pointer!!!O_O

OpenStudy (anonymous):

Anyways..........tnx for the help O'King!!Appreciate it!!

OpenStudy (kinggeorge):

No problem. Sorry about all that confusion :P

OpenStudy (anonymous):

It still shows I can bump it after 10010 mins........o_O....... NP bout dat!!

OpenStudy (anonymous):

Np on the confusion!!Happens!!Like it is now with me and the website timer!!